OpenStudy (anonymous):

Solve the equation for m: A = 2(m - w). A. m = (A + w)/2 B. m = A + w/2 C. m = 2(A + w) D. m = A/2 + w

OpenStudy (anonymous):

D

OpenStudy (anonymous):

A/2=m-w a/2 + w = m

OpenStudy (anonymous):

Multiply out the brackets \[A=2m-2w\] \[2m = A+2w\] Divide by 2 \[m = (A+2w)/2 = A/2+w\]